# Asymptotically tight Lagrangian dual of smooth nonconvex problems via improved error bound of Shapley-Folkman Lemma

### v1

@@ 4.2 Error bounds via randomized rounding
Let $A_{1},A_{2},\dots,A_{k}$ be nonempty set of $\mathbb{R}^{n}$ such that ${\mathbf{conv}}(A_{i})$ is pointed. In this subsection, we adopt the geometric perspective introduced in Lemma 4.6 to refine the classical bound of $\Phi\left(\sum_{i=1}^{k}A_{i}\right)$ stated in Corollary 1. For any $\mathbf{y}\in{\mathbf{conv}}\left(\sum_{i=1}^{k}A_{i}\right)$, Lemma 4.6 implies the existence of faces $F_{i}$ of ${\mathbf{conv}}(A_{i})$ such that $\mathbf{y}=\sum_{i=1}^{k}\mathbf{y}_{i}\text{where }\mathbf{y}_{i}\in F_{i},~\forall i\in[k].$ By definition, $\Phi\left(\sum_{i=1}^{k}A_{i}\right)$ quantifies how close $\mathbf{y}\in\sum_{i=1}^{k}{\mathbf{conv}}(A_{i})$ is to some point $\mathbf{y}^{\prime}\in\sum_{i=1}^{k}A_{i}$ in the Euclidean norm, that is, $\|\mathbf{y}-\mathbf{y}^{\prime}\|_{2}\leq\Phi(\sum_{i=1}^{k}A_{i})$. We observe that a good $\mathbf{y}^{\prime}$ can be constructed via randomly sampling some points from $F_{i}$. We first establish some results that is used to obtain such sampling.
@@
[…]
Proof 4.8.
Fix some index $j$. Write
$x=(1-\lambda_{j})\,\underbrace{\Big(\tfrac{1}{1-\lambda_{j}}\sum_{i\neq j}\lambda_{i}x_{i}\Big)}_{=:y}\;+\;\lambda_{j}x_{j}.$
By convexity of $C$, we have $y\in C$ and $x_{j}\in C$. Since $0<\lambda_{j}<1$ and $x\in F$, the point $x$ is a strict convex combination of $y$ and $x_{j}$. By the definition of a face, we conclude $y\in F$ and $x_{j}\in F$. As this holds for each $j$, all $x_{i}\in F$.
@@
Proof 4.9.
Consider any $\mathbf{x}\in F\subseteq{\mathbf{conv}}(A)$, by the definition of $\Xi(A)$, there exists $T\subseteq A$ such that $\operatorname{\operatorname{R}}(T)\leq\Xi(A)$ and $\mathbf{x}$ is a strict convex combination of some points $\mathbf{x}_{i}$ from $T$. Since $F$ is the face of ${\mathbf{conv}}(A)$, Claim 1 implies that $\mathbf{x}_{i}\subseteq F$. Since $\mathbf{x}$ is an arbitrary point in $F$ and ${\mathbf{conv}}(A\cap F)\subseteq F$, this further implies that $\Xi(A\cap F)\leq\Xi(A)$.
@@
Let $A$ be a set of $\mathbb{R}^{n}$. Then for every $\mathbf{x}\in{\mathbf{conv}}(A)$, there exists a finite distribution $\mathcal{D}$ such that $\mathcal{D}$ is supported on $A$, $\mathbb{E}_{Z\sim\mathcal{D}}(Z)=\mathbf{x}$ and $\mathbb{E}(\left\lVert Z-\mathbb{E}(Z)\right\rVert_{2}^{2})\leq\Xi^{2}(A)$ .
Proof 4.11.
For any $\mathbf{x}\in{\mathbf{conv}}(A)$, by the definition of $\Xi(A)$, there exist some $T\subseteq A$ such that $\mathbf{x}\in{\mathbf{conv}}(T)$ and $\operatorname{\operatorname{R}}(T)\leq\Xi(A)$. Since the statement of the Lemma is invariant under translation, we may translate $A$ so that $\left\lVert\mathbf{y}\right\rVert_{2}\leq\Xi(A)$ for all $\mathbf{y}\in T$. By Carathéodory’s theorem, $\mathbf{x}$ can be written as
$\mathbf{x}=\sum_{i\in[n+1]}\lambda_{i}\mathbf{p}_{i}\text{ where }\lambda\in\Delta^{n+1}\text{ and }\mathbf{p}_{i}\in T.$
We define random variable $Z\sim\mathcal{D}$ that $Z=\mathbf{p}_{i}$ with probability $\lambda_{i}$. It is clear that $\mathbb{E}_{Z\sim D}(Z)=\mathbf{x}$ and
$\mathbb{E}(\left\lVert Z-\mathbb{E}(Z)\right\rVert_{2}^{2})=\mathbb{E}(\left\lVert Z\right\rVert_{2}^{2})-\left\lVert\mathbb{E}Z\right\rVert_{2}^{2}\leq\mathbb{E}(\left\lVert Z\right\rVert_{2}^{2})\leq\Xi^{2}(A).$
where the last line uses that fact that $\left\lVert\mathbf{p}\right\rVert_{2}\leq\Xi(A)$
@@
Let $A_{1},A_{2},\dots,A_{k}$ be nonempty subsets of $\mathbb{R}^{n}$. Let $F_{i}$ be a face of ${\mathbf{conv}}(A_{i}),\forall i\in[k]$ and let $\mathcal{I}:=\{i\in[k]:\dim(F_{i})>0\}$. Then it follows that
$\Phi\left(\sum_{i=1}^{k}A_{i}\cap F_{i}\right)\leq\sqrt{\sum_{i\in\mathcal{I}}\Xi^{2}(F_{i})}\leq\sqrt{\sum_{i\in\mathcal{I}}\Xi^{2}(A_{i})}.$
Proof 4.13.
For every $\mathbf{x}\in{\mathbf{conv}}\left(\sum_{i=1}^{k}A_{i}\cap F_{i}\right)=\sum_{i=1}^{k}{\mathbf{conv}}(A_{i}\cap F_{i})$, there exits some $\mathbf{x}_{i}$ such that $\mathbf{x}=\sum_{i=1}^{k}\mathbf{x}_{i}$ and $\mathbf{x}_{i}\in{\mathbf{conv}}(A_{i}\cap F_{i})$. We aim to construct $\mathbf{y}_{i}$ such that $\mathbf{y}_{i}\in A_{i}$ and $\left\lVert\left(\sum_{i=1}^{k}\mathbf{x}_{i}\right)-\left(\sum_{i=1}^{k}\mathbf{y}_{i}\right)\right\rVert_{2}\leq\sqrt{\sum_{i\in\mathcal{I}}\Xi^{2}(F_{i})}$. For any $i\not\in\mathcal{I}$, $\dim(F_{i})=0$ and therefore it is a extreme point so that $\mathbf{x}_{i}\in A_{i}\cap F_{i}$. For $i\in I$, by Lemma 4.10, there exists a distribution $Z_{i}\sim\mathcal{D}_{i}$ such that $\mathbb{E}_{Z_{i}\sim\mathcal{D}_{i}}(Z_{i})=\mathbf{x}_{i}$ and $\mathcal{D}_{i}$ is supported on $F_{i}\cap A_{i}$ and $\mathbb{E}(\left\lVert Z_{i}-\mathbf{y}_{i}\right\rVert_{2}^{2})\leq\Xi^{2}(A_{i}\cap F_{i})\leq\beta^{2}$. Then we (randomly) construct $\mathbf{y}=\mathbf{y}_{1}+\mathbf{y}_{2}+\dots\mathbf{y}_{k}$ in the following way:
@@
$\left\lVert\left(\sum_{i=1}^{k}\mathbf{x}_{i}\right)-\left(\sum_{i=1}^{k}\mathbf{y}_{i}\right)\right\rVert_{2}\leq\sqrt{\sum_{i\in\mathcal{I}}\Xi^{2}(F_{i})}.$
Since $\mathbf{x}$ is arbitrary and by Claim 2, this implies that $\Phi\left(\sum_{i=1}^{k}A_{i}\cap F_{i}\right)\leq\sqrt{\sum_{i\in\mathcal{I}}\Xi^{2}(F_{i})}\leq\sqrt{\sum_{i\in\mathcal{I}}\Xi^{2}(A_{i})}.$
@@ 4.3 Improved error bounds via local geometry
For each $\mathbf{y}=\sum_{i=1}^{k}\mathbf{y}_{i}\in{\mathbf{conv}}\!\left(\sum_{i=1}^{k}A_{i}\right)$, Lemma 4.6 implies that each $\mathbf{y}_{i}$ lies in the Minkowski sum of certain faces $F_{i}$ of ${\mathbf{conv}}(A_{i})$. The error bound presented in Lemma 4.12 can be interpreted as follows: we attempt to round $\mathbf{y}$ as a whole to a point in $\sum_{i=1}^{k}(F_{i}\cap A_{i})$. This perspective naturally suggests a potential further refinement of Lemma 4.12, wherein we aim to round $\mathbf{y}$ directly to a point in $\sum_{i=1}^{k}A_{i}$. However, this ambitious idea encounters a significant challenge, as it requires full characterization of the nonconvex sets $A_{i}$, which is typically unavailable in practice. To address this limitation, we instead focus on exploiting only the local geometric properties of each $A_{i}$.
@@
Let $A$ be an nonempty set. For any $Q\subseteq{\mathbf{conv}}(A)$, we call $H\subseteq A$ a hidden convex component associated with $Q$ if
@@
Figure 1: Epigraph of $y=x^{4}-x^{2}$ and the star point associates a $l_{2}$ ball as a hidden convex component.
The definition above is illustrated in figure 1. Such hidden convex component may not be defined for general $Q$. On the other hand, for each extreme point $\mathbf{v}_{i}$ of ${\mathbf{conv}}(A_{i})$, there always exists a trivial hidden convex component—namely the singleton set $H_{i}=\{\mathbf{v}_{i}\}$. It is immediate that the existence of such trivial hidden convex components does not drive $\Phi\!\left(\sum_{i=1}^{k}A_{i}\right)$ to zero. For instance, take $A_{i}=\{0,1\}^{n}$ for all $i\in[k]$. Their Minkowski sum $\sum_{i=1}^{k}A_{i}$ is contained in the integer lattice $\{0,1,\dots,k\}^{n}$, and clearly $\Phi\!\left(\sum_{i=1}^{k}A_{i}\right)$ does not vanish. A natural next attempt is to require each hidden convex component $H_{i}$ to have nonempty interior. However, as illustrated in Figure 4, the mere existence of hidden convex components with non-vanishing interior still does not guarantee any improvement in the convexification of the Minkowski sum. In fact, diminishing the nonconvexity measure $\Phi\!\left(\sum_{i=1}^{k}A_{i}\right)$ requires smoothness condition of $H_{i}$.
@@
A differentiable function $f(\cdot):\mathbb{R}^{n}\to\mathbb{R}$ is called $L$-smooth if $f(\mathbf{y})\leq f(\mathbf{x})+\left\langle{\mathbf{y}-\mathbf{x}},{\nabla f(\mathbf{x})}\right\rangle+\dfrac{L}{2}\left\lVert\mathbf{y}-\mathbf{x}\right\rVert_{2}^{2},\forall\mathbf{y}\in\mathbb{R}^{n}$.
@@
Let $C\subseteq\mathbb{R}^{n}$ be a nonempty closed convex set with 0 with its interior. The gauge function associated with $C$ is
@@
For a hidden convex component $H$ with 0 in its interior, we quantify its smoothness by the smoothness of the square gauge function $\left\lVert\cdot\right\rVert_{H}^{2}$, which is related to the concept of $(2,D)$-smooth in functional analysis [molinaro2023strong].
@@
It is therefore straightforward to see that $\Phi^{K}(A)=\Phi^{K}({\mathbf{conv}}(A),A)$. Moreover, it is worth mentioning that both $\Phi(A)$ and $\Xi(A)$ are invariant under translating $A$.
@@
Let $A_{1},A_{2},\dots,A_{k}$ be nonempty subsets of $\mathbb{R}^{n}$ and $\Xi(A_{i})\leq\beta,\forall i\in[k]$. Let $F_{i}$ be a face of ${\mathbf{conv}}(A_{i}),\forall i\in[k]$ and $\mathcal{I}:=\{i\in[k]:\dim(F_{i})>0\}$. Let $H_{i}$ be a hidden convex component associated with $F_{i}$ for all $i\in[k]\setminus\mathcal{I}$ and $H=\sum\limits_{i\in[k]\setminus\mathcal{I}}H_{i}$. Suppose $H$ has nonempty interior, let $\mathcal{H}$ be a translation of $H$ that contains 0 in its interior and if we have $\left\lVert\cdot\right\rVert_{\mathcal{H}}^{2}$ is $L$-smooth, then
$\Phi^{\mathcal{H}}\left(\sum_{i=1}^{k}F_{i},\sum_{i=1}^{k}A_{i}\right)\leq\sqrt{1+\dfrac{L}{2}\sum_{i\in\mathcal{I}}^{n}\Xi^{2}(F_{i})}-1.$
Proof 4.18.
$\Phi^{\mathcal{H}}\left(\sum_{i=1}^{k}F_{i},\sum_{i=1}^{k}A_{i}\right)$ is equivalent to the maximum possible distance (under $\left\lVert\cdot\right\rVert_{\mathcal{H}}$) from points in $\sum_{i=1}^{k}F_{i}$ to $\sum_{i=1}^{k}A_{i}$. If $\sum_{i=1}^{k}F_{i}\subseteq\mathcal{F}$ for some $\mathcal{F}\subseteq\mathbb{R}^{n}$, it is clear that $\Phi^{\mathcal{H}}\left(\sum_{i=1}^{k}F_{i},\sum_{i=1}^{k}A_{i}\right)\leq\Phi^{\mathcal{H}}\left(\mathcal{F},\sum_{i=1}^{k}A_{i}\right)$. We derive an upper bound of this quantity by choosing $\mathcal{F}:=\left(\sum\limits_{i\in\mathcal{I}}F_{i}\right)+\left(\sum\limits_{i\in[k]\setminus\mathcal{I}}H_{i}\right)$. We try to show that any point in $\left(\sum\limits_{i\in\mathcal{I}}F_{i}\right)+\left(\sum\limits_{i\in[k]\setminus\mathcal{I}}H_{i}\right)$ can be rounded to some point in $\left(\sum\limits_{i\in\mathcal{I}}F_{i}\cap A_{i}\right)+\left(\sum\limits_{i\in[k]\setminus\mathcal{I}}H_{i}\right)\subseteq\sum_{i=1}^{k}A_{i}$ with a desired distance.
Since the target statement is invariant under translation, we first translate $A_{i}$ and $H_{i}$ so that $\sum\limits_{i\in[k]\setminus\mathcal{I}}H_{i}=\mathcal{H}$. Therefore, any point $\mathbf{p}$ in $\left(\sum\limits_{i\in\mathcal{I}}F_{i}\right)+\left(\sum\limits_{i\in[k]\setminus\mathcal{I}}H_{i}\right)$, can be written as $\mathbf{p}=\left(\sum\limits_{i\in\mathcal{I}}\mathbf{p}_{i}\right)+\mathbf{v}$ where $\mathbf{p}_{i}\in F_{i}$ and $\mathbf{v}\in\mathcal{H}$. By Lemma 4.10, we can equip each $F_{i}$ with a distribution $Z_{i}\sim D_{i}$ with support on $A_{i}\cap F_{i}$ such that $\mathbb{E}_{Z_{i}\sim\mathcal{D}_{i}}(Z_{i})=\mathbf{p}_{i}$ and $\mathbb{E}(\left\lVert Z_{i}-\mathbb{E}(Z_{i})\right\rVert_{2}^{2})\leq\Xi^{2}(F_{i})$. Now we try to bound
$\mathbb{E}\left(\left\lVert\left(\sum_{i\in\mathcal{I}}Z_{i}\right)-\mathbf{p}\right\rVert_{\mathcal{H}}^{2}\right)$ $=\mathbb{E}\left(\left\lVert\left(\sum_{i\in\mathcal{I}}Z_{i}-\mathbf{p}_{i}\right)-\mathbf{v}\right\rVert_{\mathcal{H}}^{2}\right)$

### v2

@@ 4.2 Error bounds via randomized rounding
Let $A_{1},A_{2},\dots,A_{k}$ be nonempty set of $\mathbb{R}^{n}$ such that ${\mathbf{conv}}(A_{i})$ is closed and pointed. For any $\mathbf{y}\in{\mathbf{conv}}\left(\sum_{i=1}^{k}A_{i}\right)$, Lemma 4.7 implies the existence of faces $F_{i}$ of ${\mathbf{conv}}(A_{i})$ such that $\mathbf{y}=\sum_{i=1}^{k}\mathbf{y}_{i}\text{ where }\mathbf{y}_{i}\in F_{i},~\forall i\in[k].$ By definition, $\Phi\left(\sum_{i=1}^{k}A_{i}\right)$ quantifies how close $\mathbf{y}\in\sum_{i=1}^{k}{\mathbf{conv}}(A_{i})$ is to some point $\mathbf{y}^{\prime}\in\sum_{i=1}^{k}A_{i}$ in the Euclidean norm, that is, $\|\mathbf{y}-\mathbf{y}^{\prime}\|_{2}\leq\Phi(\sum_{i=1}^{k}A_{i})$.
The central idea of this section is that a suitable $\mathbf{y}^{\prime}$ can be constructed by randomly sampling points from each $F_{i}$. We first establish some preliminary results that facilitate this sampling approach.
@@
Let $C$ be a closed convex set and let $F$ be a face of $C$. For any $\mathbf{x}\in F$, suppose $\mathbf{x}$ is a strict convex combination of $\mathbf{x}_{i}$s, that is, $\mathbf{x}=\sum_{i}\lambda_{i}\mathbf{x}_{i}$, with $\lambda_{i}>0$ and $\sum_{i}\lambda_{i}=1$. If $\mathbf{x}_{i}\in C$ for all $i$, then $\mathbf{x}_{i}\in F$ for all $i$.
Proof 4.10.
Directly follows from the definition of a face.
@@
Proof 4.11.
Consider any $\mathbf{x}\in F\subseteq{\mathbf{conv}}(A)$, by the definition of $\Xi(A)$, there exists $T\subseteq A$ such that $\operatorname{\operatorname{R}}(T)\leq\Xi(A)$ and $\mathbf{x}$ is a strict convex combination of some points $\mathbf{x}_{i}$ from $T$. Since $F$ is the face of ${\mathbf{conv}}(A)$, Claim 3 implies that $\mathbf{x}_{i}\subseteq F$. Since $\mathbf{x}$ is an arbitrary point in $F$ and ${\mathbf{conv}}(A\cap F)\subseteq F$, this implies that $\Xi(A\cap F)\leq\Xi(A)$.
The above claim is interesting because, in general $\Xi(A\cap B)$ may not be smaller than $\Xi(A)$ or $\Xi(B)$.
@@
Let $A\subseteq\mathbb{R}^{n}$. Then for every $\mathbf{x}\in{\mathbf{conv}}(A)$, there exists a finite distribution $\mathcal{D}$ such that $\mathcal{D}$ is supported on $A$, $\mathbb{E}_{Z\sim\mathcal{D}}(Z)=\mathbf{x}$ and $\mathbb{E}(\left\lVert Z-\mathbb{E}(Z)\right\rVert_{2}^{2})\leq\Xi^{2}(A)$ .
Proof 4.13.
For any $\mathbf{x}\in{\mathbf{conv}}(A)$, by the definition of $\Xi(A)$, there exist some $T\subseteq A$ such that $\mathbf{x}\in{\mathbf{conv}}(T)$ and $\operatorname{\operatorname{R}}(T)\leq\Xi(A)$. Since the statement of the Lemma is invariant under translation, we may translate $A$ so that $\left\lVert\mathbf{y}\right\rVert_{2}\leq\Xi(A)$ for all $\mathbf{y}\in T$. By Carathéodory’s theorem, $\mathbf{x}$ can be written a convex combination of $\mathbf{p}_{i}$ and each $\mathbf{p}_{i}\in T$, that is $\mathbf{x}=\sum_{i=1}^{n+1}\lambda_{i}\mathbf{p}_{i}\text{ where }\mathbf{p}_{i}\in T$ and $\lambda$ belongs to the $n+1$ dimensional standard simplex. We define random variable $Z\sim\mathcal{D}$ that $Z=\mathbf{p}_{i}$ with probability $\lambda_{i}$. It is clear that $\mathbb{E}_{Z\sim D}(Z)=\mathbf{x}$ and
$\mathbb{E}(\left\lVert Z-\mathbb{E}(Z)\right\rVert_{2}^{2})=\mathbb{E}(\left\lVert Z\right\rVert_{2}^{2})-\left\lVert\mathbb{E}Z\right\rVert_{2}^{2}\leq\mathbb{E}(\left\lVert Z\right\rVert_{2}^{2})\leq\Xi^{2}(A),$
where the last line uses that fact that $Z$ only has support over $T$ and $\left\lVert\mathbf{y}\right\rVert_{2}\leq\Xi(A)$ for all $\mathbf{y}\in T$.
@@
Let $A_{1},A_{2},\dots,A_{k}$ be nonempty subsets of $\mathbb{R}^{n}$. Let $F_{i}$ be a face of ${\mathbf{conv}}(A_{i})$ for all $i\in[k]$ and let $\mathcal{I}:=\{i\in[k]:\dim(F_{i})>0\}$. Then it follows that
$\Phi\left(\sum_{i=1}^{k}A_{i}\cap F_{i}\right)\leq\sqrt{\sum_{i\in\mathcal{I}}\Xi^{2}(A_{i}\cap F_{i})}\leq\sqrt{\sum_{i\in\mathcal{I}}\Xi^{2}(A_{i})}.$
Proof 4.15.
For every $\mathbf{x}\in{\mathbf{conv}}\left(\sum_{i=1}^{k}A_{i}\cap F_{i}\right)=\sum_{i=1}^{k}{\mathbf{conv}}(A_{i}\cap F_{i})$ (the equality follows from Lemma 2), there exits some $\mathbf{x}_{i}$ such that $\mathbf{x}=\sum_{i=1}^{k}\mathbf{x}_{i}$ and $\mathbf{x}_{i}\in{\mathbf{conv}}(A_{i}\cap F_{i})$. We aim to construct $\mathbf{y}_{i}$ such that $\mathbf{y}_{i}\in A_{i}\cap F_{i}$ and $\left\lVert\left(\sum_{i=1}^{k}\mathbf{x}_{i}\right)-\left(\sum_{i=1}^{k}\mathbf{y}_{i}\right)\right\rVert_{2}\leq\sqrt{\sum_{i\in\mathcal{I}}\Xi^{2}(A_{i}\cap F_{i})}$.
For any $i\not\in\mathcal{I}$, $\dim(F_{i})=0$ and therefore $\mathbf{x}_{i}$ is an extreme point of ${\mathbf{conv}}(A_{i})$, so $\mathbf{x}_{i}\in A_{i}\cap F_{i}$. For $i\in\mathcal{I}$, by Lemma 4.12, there exists a distribution $Z_{i}\sim\mathcal{D}_{i}$ such that $\mathbb{E}_{Z_{i}\sim\mathcal{D}_{i}}(Z_{i})=\mathbf{x}_{i}$ and $\mathcal{D}_{i}$ is supported on $A_{i}\cap F_{i}$ and $\mathbb{E}(\left\lVert Z_{i}-\mathbf{y}_{i}\right\rVert_{2}^{2})\leq\Xi^{2}(A_{i}\cap F_{i})$ Then we (randomly) construct $\mathbf{y}=\mathbf{y}_{1}+\mathbf{y}_{2}+\dots\mathbf{y}_{k}$ in the following way:
@@
$\left\lVert\left(\sum_{i=1}^{k}\mathbf{x}_{i}\right)-\left(\sum_{i=1}^{k}\mathbf{y}_{i}\right)\right\rVert_{2}\leq\sqrt{\sum_{i\in\mathcal{I}}\Xi^{2}(A_{i}\cap F_{i})}.$
Since $\mathbf{x}$ is arbitrary and by Claim 4, this implies that $\Phi\left(\sum_{i=1}^{k}A_{i}\cap F_{i}\right)\leq\sqrt{\sum_{i\in\mathcal{I}}\Xi^{2}(A_{i}\cap F_{i})}\leq\sqrt{\sum_{i\in\mathcal{I}}\Xi^{2}(A_{i})}.$
Applying Lemma 4.7 and Lemma 4.14 directly yields the following theorem.
Theorem 4.16.
Under conditions that each ${\mathbf{conv}}(A_{i})$ is closed and pointed for all $i\in[k]$, we have that,
$\Phi\left(\sum_{i=1}^{k}A_{i}\right)\leq\sup\left\{\sqrt{\sum_{i=1}^{k}\Xi^{2}(A_{i}\cap F_{i})}:F_{i}\text{ is a face of $A_{i}$ and $\sum_{i=1}^{k}\dim(F_{i})=n$}\right\}.$
Theorem 4.16 is a refinement on the classical error bounds, due to the fact that $\Xi(A_{i}\cap F_{i})$ can be significantly smaller than $\Xi(A_{i})$ unless $F_{i}$ is a high-dimensional face. However, due to the global dimension-sum constraints, the number of such high-dimensional faces is limited.
A more concrete illustration is as follows. Consider any full-dimensional convex body $C\in\mathbb{R}^{n}$ with $\operatorname{\operatorname{R}}(C)=R$, and define $A_{i}=C\setminus\operatorname{int}(C)$ for all $i\in[k]$. In this case, every face $F_{i}$ of ${\mathbf{conv}}(A_{i})$ that is not ${\mathbf{conv}}(A_{i})$ itself, satisfies $\Xi(A_{i}\cap F_{i})=0$ because each $F_{i}\cap A_{i}$ is a convex set. Therefore, $\Xi(A_{i}\cap F_{i})\leq R$, where $\Xi(A_{i}\cap F_{i})\neq 0$, only when $F_{i}={\mathbf{conv}}(A_{i})$ and in this case $\dim(F_{i})=n$. Now applying Theorem 4.16 implies that $\Phi\!\left(\sum_{i=1}^{k}A_{i}\right)\leq\Xi(A_{1})\leq R,$ which is significantly smaller than the $\sqrt{n}R$ bound predicted by the classical error estimates derived from the Shapley–Folkman lemma.
@@ 4.3 Improved error bounds via local geometry
For each $\mathbf{y}=\sum_{i=1}^{k}\mathbf{y}_{i}\in{\mathbf{conv}}\!\left(\sum_{i=1}^{k}A_{i}\right)$, Lemma 4.7 implies that each $\mathbf{y}_{i}$ lies in the Minkowski sum of certain faces $F_{i}$ of ${\mathbf{conv}}(A_{i})$. The error bound presented in Lemma 4.14 can be interpreted as follows: we attempt to round $\mathbf{y}$ as a whole to a point in $\sum_{i=1}^{k}(A_{i}\cap F_{i})$. This perspective naturally suggests a potential further refinement of Lemma 4.14, wherein we aim to round $\mathbf{y}$ directly to a point in $\sum_{i=1}^{k}A_{i}$. However, this appealing idea encounters a significant challenge, as it requires full characterization of the nonconvex sets $A_{i}$, which is typically unavailable in practice.
The key insight of our main result is the following: We do not need to under the global structure of the nonconvex sets $A_{i}$. It is sufficient to instead focus on exploiting only the local geometric properties of each $A_{i}$.
We next introduce a definition that plays a key role in identifying the required local geometry.
@@
Let $A$ be a nonempty set. For any $Q\subseteq{\mathbf{conv}}(A)$, we call $H\subseteq A$ is a hidden convex component associated with $Q$ if
@@
Figure 1: AA is the epigraph of the function $y=x^{4}-x^{2}$. The star point is the set $Q$. The $l_{2}$ ball contained in the epigraph is a hidden convex component.
The definition above is illustrated via an example in Figure 1.
Such hidden convex component may not exist for general $Q$. On the other hand, for each extreme point $\mathbf{v}_{i}$ of ${\mathbf{conv}}(A_{i})$, there always exists a trivial hidden convex component—namely the singleton set $H_{i}=\{\mathbf{v}_{i}\}$. It is immediate that the existence of such trivial hidden convex components does not drive $\Phi\!\left(\sum_{i=1}^{k}A_{i}\right)$ to zero. For instance, take $A_{i}=\{0,1\}^{n}$ for all $i\in[k]$. Their Minkowski sum $\sum_{i=1}^{k}A_{i}$ is contained in the integer lattice $\{0,1,\dots,k\}^{n}$, and clearly $\Phi\!\left(\sum_{i=1}^{k}A_{i}\right)$ does not vanish. A natural next attempt is to require each hidden convex component $H_{i}$ to have nonempty interior. However, as discussed in Remark 4.25 the mere existence of hidden convex components with non-vanishing interior still does not guarantee any improvement in the convexification of the Minkowski sum. In fact, diminishing the nonconvexity measure $\Phi\!\left(\sum_{i=1}^{k}A_{i}\right)$ requires an appropriate smoothness condition of the hidden convex component $H_{i}$s.
@@
[5] A differentiable function $f(\cdot):\mathbb{R}^{n}\to\mathbb{R}$ is called $L$-smooth if $f(\mathbf{y})\leq f(\mathbf{x})+\left\langle{\mathbf{y}-\mathbf{x}},{\nabla f(\mathbf{x})}\right\rangle+\dfrac{L}{2}\left\lVert\mathbf{y}-\mathbf{x}\right\rVert_{2}^{2},\forall\mathbf{y}\in\mathbb{R}^{n}$, where $\nabla f(\cdot)$ is the gradient.
@@
[21] Let $C\subseteq\mathbb{R}^{n}$ be a nonempty closed convex set with 0 with its interior. The gauge function associated with $C$ is
@@
For a hidden convex component $H$ with 0 in its interior, we quantify its smoothness by the smoothness of the square gauge function $\left\lVert\cdot\right\rVert_{H}^{2}$, which is related to the concept of $(2,D)$-smooth in functional analysis [25].
@@
It is therefore straightforward to see that $\Phi^{K}(A)=\Phi^{K}({\mathbf{conv}}(A),A)$. Moreover, we note that $\Phi^{K}(A,D)$ is invariant under translation of $A$ and $D$ by the same vector.
@@
Let $A_{1},A_{2},\dots,A_{k}$ be nonempty subsets of $\mathbb{R}^{n}$ and $\Xi(A_{i})\leq\beta,\forall i\in[k]$. Let $F_{i}$ be a face of ${\mathbf{conv}}(A_{i})$, for all $i\in[k]$ and let $\mathcal{I}:=\{i\in[k]:\dim(F_{i})>0\}$. Let $H_{i}$ be a hidden convex component associated with $F_{i}$ for all $i\in[k]\setminus\mathcal{I}$ and $H=\sum\limits_{i\in[k]\setminus\mathcal{I}}H_{i}$. Suppose $H$ has nonempty interior, let $\mathcal{H}$ be a translation of $H$ that contains 0 in its interior. If we have $\left\lVert\cdot\right\rVert_{\mathcal{H}}^{2}$ is $L$-smooth, then
$\Phi^{\mathcal{H}}\left(\sum_{i=1}^{k}F_{i},\sum_{i=1}^{k}A_{i}\right)\leq\sqrt{1+\dfrac{L}{2}\sum_{i\in\mathcal{I}}\Xi^{2}(A_{i}\cap F_{i})}-1.$
Proof 4.21.
$\Phi^{\mathcal{H}}\left(\sum_{i=1}^{k}F_{i},\sum_{i=1}^{k}A_{i}\right)$ is equivalent to the maximum possible distance (under $\left\lVert\cdot\right\rVert_{\mathcal{H}}$) from points in $\sum_{i=1}^{k}F_{i}$ to $\sum_{i=1}^{k}A_{i}$. If $\sum_{i=1}^{k}F_{i}\subseteq\mathcal{F}$ for some $\mathcal{F}\subseteq\mathbb{R}^{n}$, it is clear that $\Phi^{\mathcal{H}}\left(\sum_{i=1}^{k}F_{i},\sum_{i=1}^{k}A_{i}\right)\leq\Phi^{\mathcal{H}}\left(\mathcal{F},\sum_{i=1}^{k}A_{i}\right)$. We derive an upper bound of this quantity by choosing $\mathcal{F}:=\left(\sum\limits_{i\in\mathcal{I}}F_{i}\right)+\left(\sum\limits_{i\in[k]\setminus\mathcal{I}}H_{i}\right)$. We will show that for any point in $\left(\sum\limits_{i\in\mathcal{I}}F_{i}\right)+\left(\sum\limits_{i\in[k]\setminus\mathcal{I}}H_{i}\right)$ there exists a point in $\left(\sum\limits_{i\in\mathcal{I}}A_{i}\cap F_{i}\right)+\left(\sum\limits_{i\in[k]\setminus\mathcal{I}}H_{i}\right)\subseteq\sum_{i=1}^{k}A_{i}$ within the desired distance.
Since the target statement is invariant under equal translation of $\sum_{i=1}^{n}F_{i}$ and $\sum_{i=1}^{n}A_{i}$, we first translate $A_{i}$ so that $\sum\limits_{i\in[k]\setminus\mathcal{I}}H_{i}=\mathcal{H}$. Therefore, any point $\mathbf{p}$ in $\left(\sum\limits_{i\in\mathcal{I}}F_{i}\right)+\left(\sum\limits_{i\in[k]\setminus\mathcal{I}}H_{i}\right)$, can be written as $\mathbf{p}=\left(\sum\limits_{i\in\mathcal{I}}\mathbf{p}_{i}\right)+\mathbf{v}$ where $\mathbf{p}_{i}\in F_{i}$ and $\mathbf{v}\in\mathcal{H}$. By Lemma 4.12, we can equip each $F_{i}$ with a distribution $Z_{i}\sim D_{i}$ with support on $A_{i}\cap F_{i}$ such that $\mathbb{E}_{Z_{i}\sim\mathcal{D}_{i}}(Z_{i})=\mathbf{p}_{i}$ and $\mathbb{E}(\left\lVert Z_{i}-\mathbb{E}(Z_{i})\right\rVert_{2}^{2})\leq\Xi^{2}(A_{i}\cap F_{i})$. Now we try to bound
$\mathbb{E}\left(\left\lVert\left(\sum_{i\in\mathcal{I}}Z_{i}\right)-\mathbf{p}\right\rVert_{\mathcal{H}}^{2}\right)$ $=\mathbb{E}\left(\left\lVert\left(\sum_{i\in\mathcal{I}}Z_{i}-\mathbf{p}_{i}\right)-\mathbf{v}\right\rVert_{\mathcal{H}}^{2}\right).$